OpenStudy (anonymous):

is there an easy way to solve system of linear equation without using the matrix?

OpenStudy (anonymous):

-x-2y+4z=13 2x+y+7z=13 -7x-8y-2z=-7

OpenStudy (anonymous):

It's easier to use matrix. You can note right away if there is a solution to the system or if it has multiple solutions or a unique one.

OpenStudy (anonymous):

i have trouble solving the matrix

OpenStudy (anonymous):

Can you set up the system as a matrix?

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

What level of math are you in?

OpenStudy (anonymous):

i'm taking calc 1

OpenStudy (anonymous):

Sorry never learned how to solve systems using matrices until I started linear algebra so I don't know how to do it in a way that you will understand.

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

i do understand it but i get confused.

OpenStudy (anonymous):

There are three row operations you can do. You can add rows together and you can add scalar multiples and you can scale rows.

OpenStudy (anonymous):

You do that in order to get the matrix into echelon form and then into reduced echolon form

OpenStudy (anonymous):

ok. thanks :)